Mikayla went on a 48 mile trip to a soccer game. On the way back, due to road construction she had to drive 24 miles per hour slower. This made the trip take 1 hour longer. How fast did she drive to the soccer game?
Answer by mananth(16946)

let the returning speed be x mph
going to game speed = x+24
Distance = 48
1 hour more while returning
time return - time going = 1
48/x - 48/(x+24)=1
LCD= x ( x + 24 )
multiply by LCD
48 ( x + 24 ) - 48 x = 1 x ( x + 24 )
48 x + 1152 - 48 x = 1 X^2 + 24 x
1152 = 1 X^2 - 24 x
X^2 + 24 x -1152 = 0
x^2 +48x-24x-1152=0
x(x+48)-24(x+48)=0
(x-24)(x+48)=0
x=-48 OR 24
x =24
24 km/h while returning
while going speed = 24+24 = 48 km/h